# RiboFlow: Conditional De Novo RNA Sequence-Structure Co-Design via Synergistic Flow Matching

### v2

@@ 1 Introduction
Ribonucleic acid (RNA) is a programmable biomolecule that achieves precise molecular recognition through its dynamic three-dimensional structure (Wong et al., 2024), enabling applications in catalysis, biosensing, and therapeutic targeting (Wilson & Lilley, 2021; Cao et al., 2024). Advances in computational tools, exemplified by AlphaFold3 (Abramson et al., 2024), have revolutionized biomolecular structure prediction, while generative models (Lin et al., 2024; Zambaldi et al., 2024; Hua et al., 2024; Wang et al., 2024) now pioneer the design of de novo biomolecules with specific binding properties. RNA, with its structural flexibility at the tertiary level, ease of chemical synthesis in laboratory settings, and low immunogenicity in biological systems, stands out as a promising candidate for therapeutic drugs and biochemical reagents (Sun et al., 2020; Disney et al., 2022). However, existing methods for designing RNAs that bind specific small molecules—critical for therapeutic and diagnostic applications—face unresolved challenges at the intersection of data availability, interaction modeling, and structural validity.
Recent work has laid foundations for RNA design. Tools like RNAiFold (Dotu et al., 2014) and gRNAde (Joshi et al., 2024) generate sequences matching predefined secondary or tertiary structures, while RNA-FrameFlow (Anand et al., 2024), MMDiff (Morehead et al., 2023), and RNAFlow (Nori & Jin, 2024) focus on backbone generation. Yet, designing RNA for small-molecule targeting remains an open problem due to three gaps: (1) the inherent conformational flexibility of RNA requires the simultaneous and consistent consideration of both its structure and sequence (Falese et al., 2021); (2) existing models lack explicit conditioning on ligand geometry, limiting their ability to capture RNA-ligand binding dynamics; and (3) the scarcity of RNA-ligand structural data restricts the scalability and generalizability of data-driven approaches.
To bridge these gaps, we identify three key challenges. First, ensuring generated RNAs satisfy both binding specificity and biophysical validity requires co-designing sequence and structure in a synergistic framework. Second, modeling RNA’s structural flexibility—particularly its torsion angles and backbone dynamics—while preserving sequence-structure compatibility remains a complex geometric and thermodynamic problem. Third, the absence of large-scale, standardized RNA-ligand interaction datasets hinders training robust generative models.
@@
Our contributions are: (i) Task Formulation: We propose a first-of-its-kind synergistic flow-matching framework for ligand-conditioned de novo RNA design. The model incorporates torsion angle and backbone frame modeling, enabling sequence-structure co-design for specified ligands while offering controllable ligand-binding specificity. (ii) Dataset: We present RiboBind, the largest standardized RNA-ligand interaction benchmark, enabling data-driven RNA discovery. (iii) Evaluation: We develop a multi-faceted pipeline assessing structural validity and binding affinity (via docking and scoring). Experimental results demonstrate that RiboFlow outperforms state-of-the-art RNA design methods by a large margin (e.g., achieving a 2.2-fold improvement in the AF3 binding metric and a 50% increase in validity), but also showcases controllable capabilities for achieving high binding affinity to target ligands. We anticipate this work advancing RNA design toward structure-aware, ligand-conditioned design, with promising applications in therapeutics and synthetic biology.
@@ 2.1 RNA Design
Recently, the rise of deep learning has greatly advanced biomolecular design, leading to extensive attention in RNA design. Currently, RNA design can be broadly categorized into two main approaches: sequence-based and structure-based methods. Sequence-based design primarily aims to address the RNA inverse folding problem, which involves designing an RNA sequence that folds into a desired RNA structure. While early efforts (Dotu et al., 2014; Yang et al., 2017; Churkin et al., 2018; Runge et al., 2019) largely focus on RNA secondary structure information, recent studies (Tan et al., 2024; Joshi et al., 2024; Huang et al., 2024; Wong et al., 2024) begin to explore the use of RNA 3D structural information to guide sequence design. On the other hand, structure-based design, still in its early stages, encompasses approaches such as RNA backbone design (Anand et al., 2024), which enables the creation of RNA with specific structures. Another approach is RNA-protein co-design (Morehead et al., 2023; Nori & Jin, 2024), which focuses on designing RNA and protein components by accounting for their interactions. However, these methods are not capable of designing ligand-targeting RNAs, limiting their applications in broader diagnostic and synthetic biology scenarios.
@@ 2.2 Flow Matching
Flow matching (Lipman et al., 2023; Liu et al., 2023), an emerging generative modeling approach, is increasingly recognized as a compelling alternative to traditional generative models (Kingma, 2014; Goodfellow et al., 2014; Ho et al., 2020).It has extensive applications in both the computer vision and natural language processing fields (Dao et al., 2023; Gat et al., 2024; Esser et al., 2024). Within the realm of biomolecular design, the focus of this paper, researchers are actively exploring the application of flow matching for diverse molecular design tasks. For example, several studies (Zhang et al., 2024; Yim et al., 2023; Bose et al., 2024) utilize SE(3)-equivariant flow models to design small molecules and proteins, while others (Huguet et al., 2024; Campbell et al., 2024) integrate protein sequence for protein design, or leverage multiple sequence alignment (MSA) coevolutionary information (Hua et al., 2024; Wang et al., 2024) to design enzymes and antibodies. In contrast to these methods, RiboFlow integrates RNA-specific structural priors with sequence-structure consistency through synergistic flow matching, thereby deriving the benefits of both data-efficient representations for RNA inputs.
Figure 1: RNA backbone parameterization. Using the $C4^{\prime}$ atom as the origin, a local coordinate system is established using $v_{1}$ and $v_{2}$ defined by vectors along the $C4^{\prime}$-$O4^{\prime}$ and $C4^{\prime}$-$C3^{\prime}$ bonds via the Gram-Schmidt process. The remaining atoms are then parameterized by 8 torsion angles $\Phi$.
@@ 3 Preliminary
RNA Backbone Parameterization. As illustrated in Figure 1, an RNA has a backbone made of alternating phosphate groups ($P$, $OP1$, $OP2$, $O5^{\prime}$) and the sugar ribose ($C1^{\prime}$ $C5^{\prime}$, $O2^{\prime}$, $O3^{\prime}$, $O4^{\prime}$). A nitrogen atom, located at the base attachment site ($N9$ of purines or $N1$ of pyrimidines), is often included in modeling. Utilizing recent RNA parameterization methods (Morehead et al., 2023; Anand et al., 2024), we construct a local coordinate system for each nucleotide based on the $C4’$ ’, $C3’$ ’, and $O4’$ ’ atoms of the ribose. The positions of the remaining atoms are then determined by a set of eight torsion angles, $\Phi=\{\phi_{i}\}_{i=1}^{8}$, where $\phi_{i}\in\mathbb{R}^{2}$, along with bond lengths and bond angles (Gelbin et al., 1996). This parameterized approach can effectively captures the conformational flexibility of RNA.
[…]
@@ 4 Methods
In this section, we present RiboFlow, a flow matching framework for de novo RNA design conditioned on target small molecules. RiboFlow operates in two key stages: initially, it learns a distribution of structurally plausible RNAs during a synergistic co-design pre-training stage, establishing priors for molecule-conditioned generation. Subsequently, it is trained to generate high-affinity RNAs given a specific target molecule. For clarity, a comprehensive overview of the flow matching is discussed in the Appendix C, with mathematical tools and proofs used in our methodology. The architecture of RiboFlow is illustrated in Figure 2.
Figure 2: Framework of RiboFlow. We start by collecting and constructing RiboBind, a large-scale RNA-ligand interaction dataset, from the RCSB PDB database. RiboFlow is then pretrained on RNAsolo using a sequence-structure synergistic co-design strategy to enhance its geometric awareness. Next, the model is fine-tuned with RiboBind, enabling it to design RNAs with high affinity under ligand constraints. During inference, RiboFlow generates structurally valid and high-affinity RNAs tailored to different ligands. In the figure, “X” represents uncertain nucleotides.
@@ 4.1 Overview
As mentioned in the preliminaries, the $i$-th nucleotide in an RNA can be parameterized as $T^{(i)}=\{x^{(i)},r^{(i)},c^{(i)}\}$. We use time $t$ = 1 to represent the target data (i.e., the real RNA data $\mathcal{T}_{1}$), and $t$ = 0 to represent noise data (i.e., $\mathcal{T}_{0}$). Inspired by recent work of (Yim et al., 2023; Campbell et al., 2024), the conditional flow $p_{t}(\mathcal{T}_{t}|\mathcal{T}_{1})$ in the RNA generation process for a time step $t\in[0,1]$ can be expressed as:
@@ 4.2 RiboFlow on SE(3)
[…]
@@
Here, $\exp(\cdot)$ and $\log(\cdot)$ are the exponential and logarithmic maps respectively, enabling movement along the curved SE(3) manifold. Since the SE(3) can be decomposed into independent translations and rotations, we can also obtain the closed-form interpolations (Yim et al., 2023) for $\mathbb{R}^{3}$ and SO(3) separately:
@@
where $x_{0}$ sampled from $\mathcal{N}(0,\mathbf{I})$, and $r_{0}$ sampled uniformly from the rotation group SO(3), denoted as $\mathcal{U}(SO(3))$. Based on these interpolations, we can derive the conditional vector fields $u_{t}$ for the translation and rotation components respectively (Tong et al., 2024) due to their simple nature:
@@
$\small\mathcal{L}_{\mathrm{trans}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}(0,1),p_{1}({x}_{1}),\\ p_{0}({x}_{0}),p_{t}(x_{t}|x_{0},x_{1})\end{subarray}}\sum_{i=1}^{N_{t}}\left\|\bm{v_{\theta}}^{(i)}(x_{t},t)-x_{1}^{(i)}+x_{0}^{(i)}\right\|_{\mathbb{R}^{3}}^{2},$ (7)
@@ 4.3 RiboFlow on Torsion Angles
The conformational flexibility of the RNA backbone is largely determined by its 8 torsion angles. These angles, $\phi\in\Phi$, which are elements of $\mathbb{R}^{2}$, require appropriate constraints to generate physically plausible RNA structures.
Existing studies have explored modeling torsion angles on the torus (Lin et al., 2024) and demonstrated promising experimental results. However, such an approach may increase computational complexity when regressing conditional vector fields of torsion angles. Hence, we introduce a torsion angle prediction module, denoted as $\bm{v_{\phi}}(\cdot)$. This module employs a shallow ResNet architecture (He et al., 2016), which is first implemented in the structure prediction module of AlphaFold2 (Jumper et al., 2021) and has been widely adopted in recent works. By utilizing the predicted structural frames $\hat{F}_{1}=\{(\hat{x}^{(i)}_{1},\hat{r}^{(i)}_{1})\}_{i=1}^{N_{t}}$, the module can predict the ground-truth torsion angles $\Phi_{1}$ directly. The process can be formulated as follows:
@@ 4.4 RiboFlow on Nucleotide Type
Following the work of (Campbell et al., 2024), we employ a conditional flow to linearly interpolate from a uniform prior to $x_{1}^{(i)}$. This requires that the probability vector satisfies ${s}_{1}=\textrm{onehot}(c_{i})$ and ${s}_{0}=(\frac{1}{4},\dots,\frac{1}{4})$. The conditional flow of the probability vector is then given by:
@@ 4.5 Distance-Aware Ligand Guided RNA Generation
Given a ligand =$\mathcal{M}=\{(a^{(j)},b^{(j)})\}_{j=1}^{N_{m}}$, a two-stage distance-aware structure refinement process is designed to implicitly optimize the binding free energy by modeling 3D RNA-ligand interactions. In the first stage, a multi-layer perceptron (MLP) is employed to learn the embedding $h_{a}$ for atoms in the ligand. Subsequently, an invariant point attention (IPA) module predicts the initial RNA structural frame $\hat{F}_{1}=\{(\hat{x}_{1}^{(i)},\hat{r}_{1}^{(i)})\}_{i=1}^{N_{t}}$ without considering ligand interaction. Next, We compute the spatial interaction feature $h_{b}$ between the $i$-th RNA backbone atom ($i\in[1,N_{t}]$) and the $j$-th ligand atom ($j\in[1,N_{m}]$) as:
@@
$\small\mathcal{L}_{\mathrm{trans}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}(0,1),p({x}_{1}),\\ p_{0}({x}_{0}),p_{t}(x_{t}|x_{0},x_{1})\end{subarray}}\|v_{\theta}(x;\mathcal{C})-x_{1}^{(i)}+x_{0}^{(i)}\|_{\mathbb{R}^{3}}^{2}.$
@@
Algorithm 1 RiboFlow: Inference
  Input: Ligand structure $\mathcal{M}$, Sampling steps $T$, Initial RNA structure $\mathcal{T}_{0}$ with length $N$, Trained model $v_{\theta}$
  Initialize: $\textit{steps}\leftarrow 0$, $\textit{t}\leftarrow 0$, $\Delta t\leftarrow 1/T$
  Initialize: complex $\mathcal{C}_{0}$ with RNA $\mathcal{T}_{0}$ and ligand $\mathcal{M}$
  while $\textit{steps}<T$ do
     $x_{t+\Delta t}^{(i)}\leftarrow x_{t}^{(i)}+\bm{v}_{\theta}(x_{t}^{(i)},t;\mathcal{C}_{t})\Delta t$
     $r_{t+\Delta t}^{(i)}\leftarrow r_{t}^{(i)}\exp\left(\bm{v}_{\theta}(r_{t}^{(i)},t;\mathcal{C}_{t})\Delta t\right)$
     $s_{t+\Delta t}^{(i)}\leftarrow\textrm{norm}\big{(}s_{t}^{(i)}+\bm{v}_{\theta}(s_{t}^{(i)},t;\mathcal{C}_{t})\Delta t\big{)}$
     Sample nucleotide type: $c_{t+\Delta t}^{(i)}\sim s_{t+\Delta t}^{(i)}$
     $t\leftarrow t+\Delta t$, $\textit{steps}\leftarrow\textit{steps}+1$
  end while
  Calculate: $\hat{\Phi}_{1}^{(i)}\leftarrow\bm{v}_{\phi}(\hat{x}_{1}^{(i)},\hat{r}_{1}^{(i)})$
  Return: Final complex $\mathcal{C}_{1}$

### v3

# RiboFlow: Conditional De Novo RNA Co-Design via Synergistic Flow Matching

@@ 1 Introduction
Ribonucleic acid (RNA) is a programmable biomolecule that achieves precise molecular recognition through its dynamic three-dimensional structure wong2024deep , enabling applications in catalysis, biosensing, and therapeutic targeting wilson2021potential ; cao2024identification . Advances in computational tools, exemplified by AlphaFold3 abramson2024accurate , have revolutionized biomolecular structure prediction, while generative models linppflow ; zambaldi2024novo ; hua2024enzymeflow ; wang2024retrieval now pioneer the design of de novo biomolecules with specific binding properties. RNA, with its structural flexibility at the tertiary level, ease of chemical synthesis in laboratory settings, and low immunogenicity in biological systems, stands out as a promising candidate for therapeutic drugs and biochemical reagents sun2020rldock ; childs2022targeting . However, existing methods for designing RNAs that bind specific small molecules—critical for therapeutic and diagnostic applications—face unresolved challenges at the intersection of data availability, interaction modeling, and structural validity.
Recent work has laid foundations for RNA design. Tools like RNAiFold dotu2014complete and gRNAde joshi2024grnade generate sequences matching predefined secondary or tertiary structures, while RNA-FrameFlow anand2024rnaframeflow , MMDiff morehead2023towards , and RNAFlow nori2024rnaflow focus on backbone generation. Yet, designing RNA for small-molecule targeting remains an open problem due to three gaps: (1) the inherent conformational flexibility of RNA requires the simultaneous and consistent consideration of both its structure and sequence falese2021targeting ; (2) existing models lack explicit conditioning on ligand geometry, limiting their ability to capture RNA-ligand binding dynamics; and (3) the scarcity of RNA-ligand structural data restricts the scalability and generalizability of data-driven approaches.
To bridge these gaps, we identify three key challenges. First, ensuring generated RNAs satisfy both binding specificity and biophysical validity requires co-designing sequence and structure in a synergistic framework. Second, modeling the structural flexibility of RNA, especially its torsion angles and backbone dynamics, while maintaining sequence-structure compatibility, remains a complex problem involving both geometric and thermodynamic considerations. Third, the absence of large-scale, standardized RNA-ligand interaction datasets hinders training robust generative models.
@@
Our contributions are: (i) Task Formulation: We propose a first-of-its-kind synergistic flow-matching framework for ligand-conditioned de novo RNA design. The model incorporates torsion angle and backbone frame modeling, enabling sequence-structure co-design for specified ligands while offering controllable ligand-binding specificity. (ii) Dataset: We present RiboBind, a large standardized RNA-ligand interaction benchmark, enabling data-driven RNA discovery. (iii) Evaluation: We develop a multi-faceted pipeline assessing structural validity and binding affinity (via docking and scoring). Experimental results demonstrate that RiboFlow outperforms state-of-the-art RNA design methods by a large margin (e.g., achieving a 2.2-fold improvement in the AF3 binding metric and a 50% increase in validity), but also showcases controllable capabilities for achieving high binding affinity to target ligands. We anticipate this work advancing RNA design toward structure-aware, ligand-conditioned design, with promising applications in therapeutics and synthetic biology.
@@ 2.1 RNA Design
Currently, RNA design can be broadly categorized into two main approaches: sequence-based and structure-based methods. Sequence-based design primarily aims to address the RNA inverse folding problem, which involves designing an RNA sequence that folds into a desired RNA structure. While early efforts dotu2014complete ; yang2017rna ; churkin2018design ; rungelearning largely focus on RNA secondary structure information, recent studies tan2024rdesign ; joshi2024grnade ; huang2024ribodiffusion ; wong2024deep begin to explore the use of RNA 3D structural information to guide sequence design. On the other hand, structure-based design, still in its early stages, encompasses approaches such as RNA backbone design anand2024rnaframeflow , which enables the creation of RNA with specific structures. Another approach is RNA-protein co-design morehead2023towards ; nori2024rnaflow , which focuses on designing RNA and protein components by accounting for their interactions. However, these methods are not capable of designing ligand-targeting RNAs, limiting their applications in broader diagnostic and synthetic biology scenarios.
@@ 2.2 Flow Matching
Flow matching lipmanflow ; liuflow , an emerging generative modeling approach, is increasingly recognized as a compelling alternative to traditional generative models kingma2013auto ; goodfellow2014generative ; ho2020denoising . It has extensive applications in both the computer vision and natural language processing fields dao2023flow ; gat2024discrete ; esser2024scaling . Within the realm of biomolecular design, the focus of this paper, researchers are actively exploring the application of flow matching for diverse molecular design tasks. For example, several studies zhangflexsbdd ; yim2023fast ; bose2024se3 utilize SE(3)-equivariant flow models to design small molecules and proteins, while others huguet2024sequence ; campbell2024generative integrate protein sequence for protein design, or leverage multiple sequence alignment (MSA) coevolutionary information hua2024enzymeflow ; wang2024retrieval to design enzymes and antibodies. In contrast to these methods, RiboFlow integrates RNA-specific structural priors with sequence-structure consistency through synergistic flow matching, thereby deriving the benefits of both data-efficient representations for RNA inputs.
@@ 3 Preliminary
RNA Backbone Parameterization. As illustrated in Figure 1, an RNA has a backbone made of alternating phosphate groups ($P$, $OP1$, $OP2$, $O5^{\prime}$) and the sugar ribose ($C1^{\prime}$ - $C5^{\prime}$, $O2^{\prime}$, $O3^{\prime}$, $O4^{\prime}$). A nitrogen atom, located at the base attachment site ($N9$ of purines or $N1$ of pyrimidines), is often included in modeling. Utilizing recent RNA parameterization methods morehead2023towards ; anand2024rnaframeflow , we construct a local coordinate system for each nucleotide based on the $C4$’, $C3$’, and $O4$’ atoms of the ribose. The positions of the remaining atoms are then determined by a set of eight torsion angles, $\Phi=\{\phi_{i}\}_{i=1}^{8}$, where $\phi_{i}\in\mathbb{R}^{2}$, along with bond lengths and bond angles gelbin1996geometric . This parameterized approach can effectively captures the conformational flexibility of RNA.
Figure 1: RNA backbone parameterization. Using the $C4^{\prime}$ atom as the origin, a local coordinate system is established using $v_{1}$ and $v_{2}$ defined by vectors along the $C4^{\prime}$-$O4^{\prime}$ and $C4^{\prime}$-$C3^{\prime}$ bonds via the Gram-Schmidt process.
Notations and Problem Formulation. We consider a binding system comprising an RNA-ligand pair $\mathcal{C}=\{\mathcal{T},\mathcal{M}\}$, where the RNA, denoted by $\mathcal{T}$, consists of $N_{t}$ nucleotides, and the ligand, denoted by $\mathcal{M}$, consists of $N_{m}$ atoms. The 3D geometry of the ligand is represented as $\mathcal{M}=\{(a^{(i)},b^{(i)})\}_{i=1}^{N_{m}}$, where $a^{(i)}\in\mathbb{R}^{n_{m}}$ indicate the atom type ($n_{m}$ is the total number of possible atom types), and $b^{(i)}\in\mathbb{R}^{3}$ denotes its 3D Cartesian coordinates. The RNA is represented by the backbone atoms of each nucleotide, and is parameterized as a set of nucleotide blocks $T^{(i)}=\{x^{(i)},r^{(i)},c^{(i)}\}$. Here, $x^{(i)}\in\mathbb{R}^{3}$ is the position of the $C4^{\prime}$ atom of the $i$-th nucleotide, $r^{(i)}\in\textrm{SO(3)}$ is a rotation matrix defining the orientation of the local frame formed by $C3^{\prime}-C4^{\prime}-O4^{\prime}$ relative to a global reference frame, and $c^{(i)}\in\{\textrm{A, C, G, U}\}$ represents the nucleotide type. The complete RNA structure is represented by $\mathcal{T}=\{T^{(i)}\}_{i=1}^{N_{t}}$. Our goal is to develop a probabilistic model that learns the conditional distribution $p(\mathcal{T}|\mathcal{M})$, i.e., generating an RNA $\mathcal{T}$ conditioned on a target ligand structure $\mathcal{M}$.
@@ 4 Methods
In this section, we present RiboFlow, a flow matching framework for de novo RNA design conditioned on target small molecules. RiboFlow operates in two key stages: initially, it learns a distribution of structurally plausible RNAs during a synergistic co-design pre-training stage, establishing priors for molecule-conditioned generation. Subsequently, it is trained to generate high-affinity RNAs given a specific target molecule. For clarity, a comprehensive overview of the flow matching is discussed in the Appendix 10, with mathematical tools and proofs used in our methodology. The architecture of RiboFlow is illustrated in Figure 2.
@@ 4.1 Overview
As mentioned in the preliminaries, the $i$-th nucleotide in an RNA can be parameterized as $T^{(i)}=\{x^{(i)},r^{(i)},c^{(i)}\}$. We use time $t$ = 1 to represent the target data (i.e., the real RNA data $\mathcal{T}_{1}$), and $t$ = 0 to represent noise data (i.e., $\mathcal{T}_{0}$). Inspired by recent work of yim2023fast ; campbell2024generative , the conditional flow $p_{t}(\mathcal{T}_{t}|\mathcal{T}_{1})$ in the RNA generation process for a time step $t\in[0,1]$ can be expressed as:
@@ 4.2 RiboFlow on SE(3)
We model RNA structure in SE(3) space using a set of structure frames, denoted as $\mathbf{F}=\{(x^{(i)},r^{(i)})\}_{i=1}^{N_{t}}$, where $x^{(i)}\in\mathbb{R}^{3}$ represents the translation and $r^{(i)}\in\textrm{SO(3)}$ represents the rotation of the $i$-th nucleotide. We define a forward process that transforms an initial noisy frame set ${F_{0}}\sim p_{0}(F_{0})$ to a target structure frame set ${F_{1}}\sim p_{1}(F_{1})$. A continuous flow $F_{t}$ between $F_{0}$ and $F_{1}$ is constructed by interpolating on the SE(3) manifold:
@@
Here, $\exp(\cdot)$ and $\log(\cdot)$ are the exponential and logarithmic maps respectively, enabling movement along the curved SE(3) manifold. Since the SE(3) can be decomposed into independent translations and rotations, we can also obtain the closed-form interpolations yim2023fast for $\mathbb{R}^{3}$ and SO(3) separately:
@@
where $x_{0}$ sampled from $\mathcal{N}(0,\mathbf{I})$, and $r_{0}$ sampled uniformly from the rotation group SO(3), denoted as $\mathcal{U}(SO(3))$. Based on these interpolations, we can derive the conditional vector fields $u_{t}$ for the translation and rotation components respectively tongimproving due to their simple nature:
@@
$\mathcal{L}_{\mathrm{trans}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}(0,1),p_{1}({x}_{1}),\\ p_{0}({x}_{0}),p_{t}(x_{t}|x_{0},x_{1})\end{subarray}}\sum_{i=1}^{N_{t}}\left\|\bm{v_{\theta}}^{(i)}(x_{t},t)-x_{1}^{(i)}+x_{0}^{(i)}\right\|_{\mathbb{R}^{3}}^{2},$ (7)
@@
Figure 2: Framework of RiboFlow. We start by collecting and constructing RiboBind, a large-scale RNA-ligand interaction dataset, from the RCSB PDB database. RiboFlow is then pretrained on RNAsolo using a sequence-structure synergistic co-design strategy to enhance its geometric awareness. Next, the model is fine-tuned with RiboBind, enabling it to design RNAs with high predicted affinity under ligand constraints. During inference, RiboFlow generates structurally valid and high-affinity RNAs tailored to different ligands. In the figure, “X” represents uncertain nucleotides.
@@ 4.3 RiboFlow on Torsion Angles
The conformational flexibility of the RNA backbone is largely determined by its 8 torsion angles. These angles, $\phi\in\Phi$, which are elements of $\mathbb{R}^{2}$, require appropriate constraints to generate physically plausible RNA structures. Existing studies have explored modeling torsion angles on the torus linppflow and demonstrated promising experimental results. However, such an approach may increase computational complexity when regressing conditional vector fields of torsion angles. Insipired by Anand et al. anand2024rnaframeflow , we introduce a torsion angle prediction module, denoted as $\bm{v_{\phi}}(\cdot)$. This module employs a shallow ResNet architecture he2016deep , which is first implemented in the structure prediction module of AlphaFold2 jumper2021highly and has been widely adopted in recent works. By utilizing the predicted structural frames $\hat{F}_{1}=\{(\hat{x}^{(i)}_{1},\hat{r}^{(i)}_{1})\}_{i=1}^{N_{t}}$, the module can predict the ground-truth torsion angles $\Phi_{1}$ directly. The process can be formulated as follows:
@@ 4.4 RiboFlow on Nucleotide Type
Following the work of campbell2024generative , we employ a conditional flow to linearly interpolate from a uniform prior to $x_{1}^{(i)}$. This requires that the probability vector satisfies ${s}_{1}=\textrm{onehot}(c_{i})$ and ${s}_{0}=(\frac{1}{4},\dots,\frac{1}{4})$. The conditional flow of the probability vector is then given by:
@@ 4.5 Distance-Aware Ligand Guided RNA Generation
Given a ligand $\mathcal{M}=\{(a^{(j)},b^{(j)})\}_{j=1}^{N_{m}}$, a two-stage distance-aware structure refinement process is designed to implicitly optimize the binding free energy by modeling 3D RNA-ligand interactions. In the first stage, a multi-layer perceptron is employed to learn the embedding $h_{a}$ for atoms in the ligand. Subsequently, an invariant point attention (IPA) module predicts the initial RNA structural frame $\hat{F}_{1}=\{(\hat{x}_{1}^{(i)},\hat{r}_{1}^{(i)})\}_{i=1}^{N_{t}}$ without considering ligand interaction. Next, We compute the spatial interaction feature $h_{b}$ between the $i$-th RNA backbone atom ($i\in[1,N_{t}]$) and the $j$-th ligand atom ($j\in[1,N_{m}]$) as:
@@
$\mathcal{L}_{\mathrm{trans}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}(0,1),p({x}_{1}),\\ p_{0}({x}_{0}),p_{t}(x_{t}|x_{0},x_{1})\end{subarray}}\|v_{\theta}(x;\mathcal{C})-x_{1}^{(i)}+x_{0}^{(i)}\|_{\mathbb{R}^{3}}^{2}.$ (17)